Database last updated on Mon, 26 Jan 2026 08:40:49 +0000 by localhost
Instrument constant changes for the week before 2025327 (2 sec, 0 rows)
SL burned out? (AP results when SL is ON) (0 sec, 3 rows)
Non-critical warnings from b-files' comments for days between 2025327 and 2025327 (0 sec, 2 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| rl: | 2025 326 06 15 17 Reset Warm -> Tepid from 2025 326 06 15 11 until 2025 326 06 15 12 . Counts: 1,0 | 1 | 3 | B32725.203 | CHURC | #203 | 06:08:32 | | 1 | -1 |
| rl: | 2025 327 09 18 22 New value set for watermark SUPPLY.VOLTAGE.HIGH | 2 | 3 | B32725.203 | CHURC | #203 | 06:08:35 | | 1 | -1 |
Measurements and tests summary.
| _jday | _loc | _serial | _skc | dt | ed | fz | hg | hp | rs | sl | uv | zs |
| 327 | CHURCHILL | #203 | ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.02201) | 2 | 3 | 33 | 12 | 12 | 2 | 8 | 21 | 36 |
(0 sec)
Relative and absolute humidity for day 2025327 (2 sec, 1 rows)
| Serial | RH, % | AH, g/m3 | 7-day RH history | Des-nt replcd days ago |
| 203 | 6.9 | 1.1 |